PITMinutes 02-15-2016

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, S. Cooper, E. Daly, K. Davis, M. Drury, C. Hovater, R. Geng, G. Krafft, C. Mounts, C. Reece, T. Reilly, R. Rimmer, T. Plawski, T. Powers, H. Zhang

Agenda and Notes

Particulate Meeting two weeks

1. Operations Status and Summary

Beam delivery to users.

Heat Measurement Status

Energy Jitter 1.6 e-4 main effect at 28 Hz. about 700 kV amplitude

Phase noise spur at 27 Hz in the 1427 MO.

install D Coupler and check LO

Beam studies tomorrow! 10:30-3:00 RF Optimization. Firmware changes? Arne requests cavity by cavity if do it.

Q0 measurements on the LERF?

Piezo work started on R100. Should be possible to get some microphonic damping. Stay tuned, code/software problems.

Regular beam studies on Tuesdays and thursdays for RF recovery.

One hour machine ops devoted to Fast X-ray detection of FE, able to tune for reduced losses in . Establish baselines for future comparisons.

C100 tests (determine best pressure "defined by highest operating gradient" to operate at)

Got 70% of cavities back up to E_maxop values, but were having cooling/instability problems when operating at high values.

destinguish/study SEL GDR trip differences (Rama)

Future mechanical investigations to elliminate vibrations in the first place may be appropriate going forward.

2. Superinsulation Damage? (???)

Cabling damage is starting to be observed.

Monitor heat load changes, especially in C100s, to see if changes in material show up here. TLDs in the tunnel (end cans and warm girders)

3. R&D for improving and maintaining CEBAF gradient (in situ?)

Improvement of the crashed cryomodules, including a C50. Is there any way to recover them. Plasma processing? SNS information: plasma processing Hydrocarbon removal. Try out on removed C75-1 would be nice. Previous experience at JLAB not good.

Microphonics in C75 worse because higher Q. Worry here!

Action Items